OpenStudy (anonymous):

If a,b are the roots of x^2-3x+1=0, find the equation whose roots are 1+a+a^2 & 1+b+b^2 . . ANSWER=x^2-12x+16=0

OpenStudy (anonymous):

sum = 3 product = 1

OpenStudy (anonymous):

sum of roots = a^2+a+1+1+b+b^2

hartnn (hartnn):

so you know a+b=3 ab=1

OpenStudy (anonymous):

hmm :)

hartnn (hartnn):

you just need a^2 6b^2, right ?

hartnn (hartnn):

a^2+b^2

OpenStudy (anonymous):

how to sum those

hartnn (hartnn):

a^2+b^2 = (a+b)^2 -2ab i think we did this last time ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2+(a+b)^2-2ab+(a+b)

hartnn (hartnn):

okk...for this, so now you have sum, just need product, ... (a^2+a+1)(1+b+b^2 )

OpenStudy (anonymous):

is this right?

hartnn (hartnn):

yes, just plug in values.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12

hartnn (hartnn):

yes.

OpenStudy (anonymous):

now product

OpenStudy (anonymous):

1+b+b^2+a+ab+ab^2+a^2+a^2b+a^2b^2

OpenStudy (anonymous):

is this right?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

whats next?

hartnn (hartnn):

1+b+b^2+a+ab+ab^2+a^2+a^2b+a^2b^2 = 1+ (a+b) + ab + (a^2+b^2) + (ab)^2 + ab (a+b)

OpenStudy (anonymous):

this is just like my anime....

hartnn (hartnn):

which term you have doubt ?

OpenStudy (anonymous):

which anime?

OpenStudy (anonymous):

16

hartnn (hartnn):

yeah, so u got the equation ?